# A continuous boostlet transform for acoustic waves in space–time

### v1

@@ 2.3 Trace waves and phase velocities
Let us consider, without loss of generality, a propagating wave with azimuth angle $\alpha=0$ and some elevation angle $\beta$ with respect to the normal of the $x$-axis. Then, using (5), the trace wavenumber $k=k_{0}\sin\beta$, and the phase velocity along the $x$-axis reads $c_{x}=c_{0}/\sin\beta$. On the one hand, for instance, a propagating wave with $\beta=\pi/2$ is recorded at a line of receivers on the $x$-axis with a phase velocity $c_{x}=c_{0}$. Similarly, if $\beta=0$, the propagating wave has infinite phase velocity. On the other hand, an evanescent wave has an imaginary $\beta$, which results in a phase velocity always smaller than the sound speed $c_{0}$ and an exponential decay along the normal of the $x$-axis. In other words, as illustrated on the right of Fig. 2, propagating waves are inside the radiation cone, with phase velocity $c_{x}>c_{0}$, and are commonly referred to as supersonic waves. Conversely, evanescent waves are outside the radiation cone, with phase velocity $c_{x}<c_{0}$, and are commonly referred to as subsonic waves.
@@ 2.4 Encoding the dispersion relation with the Poincaré group
From a geometric perspective, we have seen in the previous sections that the dispersion relation defines the radiation cone, as shown in Fig. 2. Representing a wavefront localized in the wavenumber–frequency domain needs to respect such a geometric structure. In other words, encoding wave propagation physics into the representation system means no transformation of the signal-acquisition method can modify the dispersion relation. This is equivalent to requiring the dispersion relation to be covariant with respect to the group of transformations underlying the signal representation. As shown in Fig. 1(c), experimental evidence suggests that acoustic fields are localized into phase velocity cones and hyperbolic scales (or wavenumber–frequency dilations). These theoretical and experimental arguments suggest that the Poincaré group [^2] The Poincaré group, $\mathbb{P}(1,d-1)=\mathbb{R}^{1,d-1}\ltimes\mathbb{O}(1,d-1)$, is the semi-direct product of two subgroups: (i) translations $\mathbb{R}^{1,d-1}$, and (ii) the Lorentz group $\mathbb{O}(1,d-1)$, consisting of Lorentz (hyperbolic) boosts, in $d$-dimensional space–time. is a natural candidate, as it is often used to describe $d$-dimensional space–time in special relativity theory and preserves cone symmetries. Specifically, the orbits of the Lorentz subgroup form hyperbolae in the wavenumber–frequency domain, reproducing the structures observed in experimental data. Concretely, these boosts modify the phase velocity of a wavefield while preserving its scale. Conversely, the group of (isotropic) dilations modifies the scale of the wavefield while preserving the phase velocity. For these reasons, the boostlet construction is based on the semi-direct product of the Poincaré group with dilations [^3] This is known as the Weyl group in special relativity, $\mathbb{W}(1,d-1)=\mathbb{R}^{+}\ltimes\mathbb{P}(1,d-1)$, and it is an invariance group for photons [70]. In the face of ambiguity with the Weyl–Heisenberg group in harmonic analysis, we will use the notion of a boostlet group $\mathbb{B}$ in this paper..
@@
To illustrate the above intuition in a geometric fashion, let us apply various cell partitions to the 2D wavenumber-frequency domain, as shown in Fig. 3. The radiation cone is overlaid in orange. Plane waves, wavelets, and wave atoms partition the spectrum into rectangular cells of various areas [see Figs. 3(a), (b), and (c)]. Unless the spectrum has a high enough resolution, most of the coefficients nearest the cone boundary are inside and outside, which does not respect the dispersion relation. While curvelets and shearlets, shown in Fig. 3(c), can be defined with a conic partition, their decomposition scales (corona) do not follow the hyperbolae localized into phase velocity cones as shown in Fig. 1(c), in other words, the number of shears levels required to properly represent a wavefield varies with the dilation level, making shearlets ill-suited for representing wavefields. This approach might be useful if the wave field consists of time-harmonic waves in a band-limited frequency range. Another requirement for shearlets observed in previous experimental investigations [19] is that the shearlet cones [^4] In the jargon of cone-adapted shearlets [71], their vertical and horizontal cones correspond to the far-field and near-field acoustic cones. match the radiation cone—i.e., space–time should be sampled such that the sampling wavenumber and sampling frequency are scaled by the sound speed. The directional sub-band filters proposed by Pinto and Vetterli [58], shown in Fig. 3(e), consist of band-pass filtering transient plane waves within intervals of phase velocities. These filters are closest to boostlets due to the use of phase velocity cones. The boostlet partitioning is shown in Fig. 3(f), initially proposed in [60], and composed of hyperbolae localized in phase velocity cones and a concave-shaped scaling function (more details in Sec. 3.4). The main difference between boostlets and directional filters [58] lies in the localization of narrow-band frequency content within hyperbolic bands. Localizing narrow-band frequencies in an acoustic wave field allows for a flexible description of frequency-dependent absorption and scattering phenomena, e.g., different bands can have different weights.
@@ 3 Continuous boostlet transform
The framework to handcraft the continuous boostlet transform, which follows the same spirit of previous harmonic analysis constructions [42, 43], is presented in this section. The derivations in this paper are in 2D space–time, that is, 1D space and time, and extensions to higher spatial dimensions are the subject of future work. The admissibility condition is derived using the group structure of the transform. Lastly, Parseval’s relationship associated with the boostlet transform is shown, leading us to the definition of the boostlet scaling function. We also visualize examples of boostlet functions, the scaling function, and their associated coefficients using an experimentally measured acoustic field.
@@
$D_{a}=\begin{pmatrix}a&0\\ 0&a\end{pmatrix},\,\,\,B_{\theta}=\begin{pmatrix}\cosh\theta&-\sinh\theta\\ -\sinh\theta&\cosh\theta\end{pmatrix},$
@@
$M_{a,\theta}=D_{a}B_{\theta}=B_{\theta}D_{a}=\begin{pmatrix}a\cosh\theta&-a\sinh\theta\\ -a\sinh\theta&a\cosh\theta\end{pmatrix}.$
@@
$\hat{\psi}_{a,\theta,\tau}=e^{-2\pi i\tau^{\text{T}}\xi}\hat{\psi}(M_{a,\theta}^{\text{T}}\xi).$ (8)
@@
$\mathbb{B}=\{(a,\theta,\tau):a\in\mathbb{R}^{+},\theta\in\mathbb{R},\tau\in\mathbb{R}^{2}\}.$
@@
$f\to\mathcal{B}_{f}=\left(\langle f,\psi_{a,\theta,\tau}\rangle,\langle f,\psi_{a,\theta,\tau}^{*}\rangle\right)_{(a,\theta,\tau)\in\mathbb{B}}.$
@@
How the set of boostlet functions respects the dispersion relation (and the acoustic radiation cone) is the subject of the following sections. In particular, while the above transform can be computed for any choice of functions $\psi\in L^{2}(\mathbb{R}^{2})$, it is not clear whether the original wavefield can be recovered from the transform. To investigate this, we begin with the definition of the boostlet group, which allows us to derive an admissibility condition for $\psi$ as well as several useful properties of the transform.
@@ 3.1 Boostlet group
To better understand the properties of the continuous boostlet transform defined above, we first need a more detailed characterization of the group of transformations used to generate the boostlet system.
@@
The set $\mathbb{B}=\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}^{2}$ equipped with a product normal-$\cdot$ given by
$\left(a,\theta,\tau\right)\cdot\left(a^{\prime},\theta^{\prime},\tau^{\prime}\right)=\left(aa^{\prime},\theta+\theta^{\prime},\tau+B_{\theta}D_{a}\tau^{\prime}\right)$
forms a group with identity $(1,0,0)$.
Proof.
It can be checked that $(1,0,0)$ is the identity element. The inverse of an element $(a,\theta,\tau)\in\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}^{2}$ is given by
$(a,\theta,\tau)^{-1}=\left(\frac{1}{a},-\theta,-D_{a}^{-1}B_{\theta}^{-1}\tau\right),$
since
$(a,\theta,\tau)\cdot\left(\frac{1}{a},-\theta,-D_{a}^{-1}B_{\theta}^{-1}\tau\right)=\left(a\frac{1}{a},\theta-\theta,\tau-B_{\theta}D_{a}D_{a}^{-1}B_{\theta}^{-1}\tau\right)=(1,0,0)$
and
$\left(\frac{1}{a},-\theta,-D_{a}^{-1}B_{\theta}^{-1}\tau\right)\cdot(a,\theta,\tau)=\left(\frac{1}{a}a,-\theta+\theta,-D_{a}^{-1}B_{\theta}^{-1}\tau+B_{-\theta}D_{\frac{1}{a}}\tau\right)=(1,0,0).$
The product $\cdot$ is also associative, which is shown by
$\begin{split}((a,\theta,\tau)&\cdot(a^{\prime},\theta^{\prime},\tau^{\prime}))\cdot(a^{\prime\prime},\theta^{\prime\prime},\tau^{\prime\prime})=(aa^{\prime},\theta+\theta^{\prime},\tau+B_{\theta}D_{a}\tau^{\prime})\cdot(a^{\prime\prime},\theta^{\prime\prime},\tau^{\prime\prime})\\ &=(aa^{\prime}a^{\prime\prime},\theta+\theta^{\prime}+\theta^{\prime\prime},\tau+B_{\theta}D_{a}\tau^{\prime}+B_{\theta+\theta^{\prime}}D_{aa^{\prime}}\tau^{\prime\prime})\\ &=(a(a^{\prime}a^{\prime\prime}),\theta+(\theta^{\prime}+\theta^{\prime\prime}),\tau+B_{\theta}D_{a}(\tau^{\prime}+B_{\theta^{\prime}}D_{a^{\prime}}\tau^{\prime\prime}))\\ &=(a,\theta,\tau)\cdot(a^{\prime}a^{\prime\prime},\theta^{\prime}+\theta^{\prime\prime},\tau^{\prime}+B_{\theta^{\prime}}D_{a^{\prime}}\tau^{\prime\prime})\\ &=(a,\theta,\tau)\cdot((a^{\prime},\theta^{\prime},\tau^{\prime})\cdot(a^{\prime\prime},\theta^{\prime\prime},\tau^{\prime\prime})),\end{split}$
where we have used the fact that $B_{\theta+\theta^{\prime}}=B_{\theta}B_{\theta^{\prime}}$ which follows from basic trigonometric identities. ∎
We will refer to the group $\mathbb{B}$ with the above structure as the boostlet group. More generally, this can be thought of as the semi-direct product of the Poincaré group with isotropic dilations in space–time. At this point, we find no reason to dilate space–time anisotropically because the propagation medium considered in this study is non-dispersive. This can be compared to, for example, curvelets, shearlets, and related constructions, which have strong reasons to dilate real space parabolically. Incorporating such parabolic dilations in space may be sensible when extending the boostlet group to higher spatial dimensions. As it will later become evident, however, the action of the boost matrix $B_{\theta}$ does generate anisotropic elements in space–time.
Let us now consider the action of the boostlet group $\mathbb{B}$ on the space of wavefields, here represented with functions in $L^{2}(\mathbb{R}^{2})$.
@@
Proof.
$\sigma^{-1}(a,\theta,\tau)\psi(\varsigma)=a\psi(D_{a}B_{\theta}\varsigma+\tau).$
$\langle\sigma f,g\rangle=\langle f,\sigma^{*}g\rangle=\langle f,\sigma^{-1}g\rangle,$
$\langle\sigma f,g\rangle=\int a^{-1}f\left(B_{\theta}^{-1}D_{a}^{-1}(\varsigma-\tau)\right)g(\varsigma)\,\mathrm{d}\varsigma,$
while the last term is defined as
$\langle f,\sigma^{-1}g\rangle=\int af(\varsigma)g\left(D_{a}B_{\theta}\varsigma+\tau\right)\,\mathrm{d}\varsigma.$
$\langle\sigma f,g\rangle=\langle f,\sigma^{-1}g\rangle=\int a^{-1}f\left(B_{\theta}^{-1}D_{a}^{-1}(\varsigma^{\prime}-\tau)\right)g(\varsigma^{\prime})\,\mathrm{d}\varsigma^{\prime}\,.$
In addition to its group structure, $\mathbb{B}$, as a subset of $\mathbb{R}^{4}$, also possesses a differentiable manifold structure. Consequently, $\mathbb{B}$ is a Lie group, which, among other things, lets us define a Haar measure on the group.
@@
$\mathrm{d}\mu(a,\theta,\tau)=a^{-3}\mathrm{d}a\mathrm{d}\theta\mathrm{d}\tau.$
Proof.
$\begin{split}a^{\prime\prime}&=a^{\prime}a,\\ \theta^{\prime\prime}&=\theta^{\prime}+\theta,\\ \tau^{\prime\prime}&=\tau^{\prime}+B_{\theta^{\prime}}D_{a^{\prime}}\tau.\end{split}$
Next is to compute the determinant of the Jacobian matrix:
$\begin{vmatrix}\frac{\partial a^{\prime\prime}}{\partial a}&\frac{\partial a^{\prime\prime}}{\partial\theta}&\frac{\partial a^{\prime\prime}}{\partial\tau_{x}}&\frac{\partial a^{\prime\prime}}{\partial\tau_{y}}\\ \frac{\partial\theta^{\prime\prime}}{\partial a}&\frac{\partial\theta^{\prime\prime}}{\partial\theta}&\frac{\partial\theta^{\prime\prime}}{\partial\tau_{x}}&\frac{\partial\theta^{\prime\prime}}{\partial\tau_{y}}\\ \frac{\partial\tau_{x}^{\prime\prime}}{\partial a}&\frac{\partial\tau_{x}^{\prime\prime}}{\partial\theta}&\frac{\partial\tau_{x}^{\prime\prime}}{\partial\tau_{x}}&\frac{\partial\tau_{x}^{\prime\prime}}{\partial\tau_{y}}\\ \frac{\partial\tau_{y}^{\prime\prime}}{\partial a}&\frac{\partial\tau_{y}^{\prime\prime}}{\partial\theta}&\frac{\partial\tau_{y}^{\prime\prime}}{\partial\tau_{x}}&\frac{\partial\tau_{y}^{\prime\prime}}{\partial\tau_{y}}\end{vmatrix}=\begin{vmatrix}a^{\prime}&0&0&0\\ 0&1&0&0\\ 0&0&a^{\prime}\cosh{\theta^{\prime}}&-a^{\prime}\sinh{\theta^{\prime}}\\ 0&0&-a^{\prime}\sinh{\theta^{\prime}}&a^{\prime}\cosh{\theta^{\prime}}\end{vmatrix}={a^{\prime}}^{3}.$
@@
for all $f\in L^{2}(\mathbb{R}^{2})$, where $\mathrm{d}\mu(a,\theta,\tau)$ is the left-invariant Haar measure of $\mathbb{B}$ (see Lemma 3.3).
A rigorous characterization of these admissible boostlets is given in Theorem 1.

### v2

@@ 2.3 Trace waves and phase velocities
Let us consider, without loss of generality, a propagating wave with azimuth angle $\alpha=0$ and some elevation angle $\beta$ with respect to the normal of the $x$-axis. Then, using (6), the trace wavenumber $k=k_{0}\sin\beta$, and the phase velocity along the $x$-axis reads $c_{x}=c_{0}/\sin\beta$. On the one hand, for instance, a propagating wave with $\beta=\pi/2$ is recorded at a line of receivers on the $x$-axis with a phase velocity $c_{x}=c_{0}$. Similarly, if $\beta=0$, the propagating wave has infinite phase velocity. On the other hand, an evanescent wave has an imaginary $\beta$, which results in a phase velocity always smaller than the sound speed $c_{0}$ and an exponential decay along the normal of the $x$-axis. In other words, as illustrated on the right of Figure 2, propagating waves are inside the radiation cone, with phase velocity $c_{x}>c_{0}$, and are commonly referred to as supersonic waves. Conversely, evanescent waves are outside the radiation cone, with phase velocity $c_{x}<c_{0}$, and are commonly referred to as subsonic waves.
@@ 2.4 Encoding the dispersion relation with the Poincaré group
From a geometric perspective, we have seen in the previous sections that the dispersion relation defines the radiation cone, as shown in Figure 2. Representing a wavefront localized in the wavenumber–frequency domain needs to respect such a geometric structure. In other words, encoding wave propagation physics into the representation system means no transformation of the signal-acquisition method can modify the dispersion relation. This is equivalent to requiring the dispersion relation to be covariant with respect to the group of transformations underlying the signal representation.
@@
As experimentally shown in Figure 1(c), acoustic fields are localized into phase velocity cones and hyperbolic scales (or wavenumber–frequency dilations). These theoretical and experimental arguments suggest that the Poincaré group is a natural candidate, as it preserves cone symmetries and much of its mathematical properties have been studied in Poincaré-based wavelet systems [13]. Specifically, the orbits of the Lorentz subgroup form hyperbolae in the wavenumber–frequency domain, reproducing the structures observed in experimental data. Concretely, these boosts modify the phase velocity of a wavefield, $c_{x}$, while preserving its scale. Conversely, the group of (isotropic) dilations modifies the scale of the wavefield while preserving the phase velocity. For these reasons, the boostlet construction is based on the semi-direct product of the Poincaré group with dilations.
@@
To illustrate the above intuition in a geometric fashion, let us apply various cell partitions to the 2D wavenumber–frequency domain, as shown in Figure 3. The radiation cone is overlaid in orange. Plane waves, wavelets, and wave atoms partition the spectrum into rectangular cells of various areas (see Figures  3(a), (b), and (c)). Unless the spectrum has a high enough resolution, most of the coefficients nearest the cone boundary are inside and outside, which does not respect the dispersion relation. While curvelets and shearlets, shown in Figure 3(c), can be defined with a conic partition, their decomposition scales (corona) do not follow the hyperbolae localized into phase velocity cones as shown in Figure 1(c). The directional sub-band filters proposed by Pinto and Vetterli [28], shown in Figure 3(e), consist of band-pass filtering transient plane waves within intervals of phase velocities. These filters are closest to boostlets due to the use of phase velocity cones. The boostlet partitioning is shown in Figure 3(f), initially proposed in [66], and composed of hyperbolae localized in phase velocity cones and a concave-shaped scaling function (more details in Section 3.3). The main difference between boostlets and directional filters lies in the localization of narrow-band frequency content within hyperbolic bands. Localizing narrow-band frequencies in an acoustic wave field allows for a flexible description of frequency-dependent absorption and scattering phenomena, e.g., different bands can have different weights.
@@ 3 Continuous boostlet transform
The framework to handcraft the continuous boostlet transform, which follows the same spirit of previous harmonic analysis constructions [20, 21, 67], is presented in this section. The derivations in this paper are in 2D space–time, that is, 1D space and time, and extensions to higher spatial dimensions are the subject of future work. The admissibility condition is derived using the group structure of the transform. Lastly, Parseval’s relationship associated with the boostlet transform is shown, leading us to the definition of the boostlet scaling function for finite dilations.
@@
$D_{a}=\begin{pmatrix}a&0\\ 0&a\end{pmatrix},\,\,\,B_{\theta}=\begin{pmatrix}\cosh\theta&-\sinh\theta\\ -\sinh\theta&\cosh\theta\end{pmatrix},$ (13)
@@
$M_{a,\theta}=D_{a}B_{\theta}=B_{\theta}D_{a}=\begin{pmatrix}a\cosh\theta&-a\sinh\theta\\ -a\sinh\theta&a\cosh\theta\end{pmatrix}.$ (14)
@@
$\hat{\psi}_{a,\theta,\tau}(\xi)=a\,e^{-2\pi i\tau^{\text{T}}\xi}\hat{\psi}(M_{a,\theta}^{\text{T}}\xi).$ (17)
@@
$\mathbb{B}=\{(a,\theta,\tau):a\in\mathbb{R}^{+},\theta\in\mathbb{R},\tau\in\mathbb{R}^{2}\}.$ (19)
@@
$f\to\mathcal{B}_{f}=\left(\langle f,\psi_{a,\theta,\tau}\rangle,\langle f,\psi_{a,\theta,\tau}^{*}\rangle\right)_{(a,\theta,\tau)\in\mathbb{B}}.$ (20)
@@
How the set of boostlet functions respects the dispersion relation (and the acoustic radiation cone) is the subject of the following sections. In particular, while the above transform can be computed for any choice of functions $\psi\in L^{2}(\mathbb{R}^{2})$, it is not clear whether the original wavefield can be recovered from the transform. To investigate this, we begin with the definition of the boostlet group, which allows us to derive the admissibility condition for $\psi$ as well as several useful properties of the transform.
@@ 3.1 Boostlet group
To better understand the properties of the continuous boostlet transform defined above, we first need a more detailed characterization of the group of transformations used to generate the boostlet system. The boostlet group is also known in the wavelet literature as the affine Poincaré group, the Weyl–Poincaré group, and the similitude group [13]. The proof of the following lemma is included in A.
@@
The set $\mathbb{B}=\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}^{2}$ equipped with a product $\cdot$ given by
$\left(a,\theta,\tau\right)\cdot\left(a^{\prime},\theta^{\prime},\tau^{\prime}\right)=\left(aa^{\prime},\theta+\theta^{\prime},\tau+B_{\theta}D_{a}\tau^{\prime}\right)$ (21)
forms a group with identity (1,0,0)(1,0,0)( 1 , 0 , 0 ).
We will refer to the group $\mathbb{B}$ with the above structure as the boostlet group. More generally, this can be thought of as the semi-direct product of the Poincaré group with isotropic dilations in space–time [13]. At this point, we find no reason to dilate space–time anisotropically because the propagation medium considered in this study is non-dispersive. This can be compared to, for example, curvelets, shearlets, and related constructions, which have strong reasons to dilate real space parabolically. Incorporating such parabolic dilations in space may be sensible when extending the boostlet group to higher spatial dimensions. As it will later become evident, however, the action of the boost matrix $B_{\theta}$ does generate anisotropic elements in space–time.
Let us now consider the action of the boostlet group $\mathbb{B}$ on the space of wavefields, here represented with functions in $L^{2}(\mathbb{R}^{2})$. In addition to its group structure, $\mathbb{B}$, as a subset of $\mathbb{R}^{4}$, also possesses a differentiable manifold structure. Consequently, $\mathbb{B}$ is a Lie group, which, among other things, lets us define a Haar measure on the group. We include the proofs of the following lemmata in B and C for completeness.
@@
$\mathrm{d}\mu(a,\theta,\tau)=a^{-3}\mathrm{d}a\mathrm{d}\theta\mathrm{d}\tau.$ (23)
@@
for all $f\in L^{2}(\mathbb{R}^{2})$, where $\mathrm{d}\mu(a,\theta,\tau)$ is the left-invariant Haar measure of $\mathbb{B}$ in Eq. (23).
A rigorous characterization of these admissible boostlets is given later in Theorem 3.1.
@@ 3.2 Admissible boostlets
Let us begin with a lemma that characterizes the square integrability of the boostlet functions and their corresponding admissibility. The proof of this lemma is found in D.